# Adaptive Error Correction for Entanglement Distillation

### v1

@@ Abstract
Quantum network applications impose a variety of requirements on entanglement resources in terms of rate, fidelity, latency, and more. The repeaters in the quantum network must combine good methods for entanglement generation, effective entanglement distillation, and smart routing protocols to satisfy these application requirements. In this work, we focus on quantum error correction-based entanglement distillation in a linear chain of quantum repeaters. While conventional approaches reuse the same distillation scheme over multiple hop lengths after entanglement swaps, we propose a novel adaptive error correction scheme that boosts end-to-end metrics. Specifically, depending on the network operating point, we adapt the code used in distillation over successive rounds to monotonically increase the rate while also improving fidelity. We demonstrate the effectiveness of this strategy using three codes: $\llbracket 9,1,3\rrbracket$, $\llbracket 9,2,3\rrbracket$, $\llbracket 9,3,3\rrbracket$. We compare the performance of four different protocols that combine the codes in different ways, where we define a new performance metric, efficiency, that incorporates both overall rate and fidelity. While we highlight our innovation under minimal assumptions on noise, the method can be easily generalized to realistic network settings. By combining our approach with good entanglement generation methods and smart routing protocols, we can achieve application requirements in a systematic, resource-efficient, way.
Index Terms:
Quantum repeaters, entanglement distillation, stabilizer code, depolarizing channel, Werner state
@@ I Introduction
Quantum networks fundamentally differ from classical networks since they rely upon entanglement distribution across the network nodes. Since direct transmission of quantum information across long distances suffers from an exponential scaling of loss [1], quantum repeaters are critical for wide area quantum networks. A repeater generates entanglement across single hops – short links with its neighboring nodes, which could be end-users or other repeaters – and performs entanglement swaps to create longer-range entanglement across longer hops. Since each link introduces loss, and swaps involve local operations that are noisy, entanglement distillation is used to distill fewer high fidelity entangled states from many noisy entangled states. Such distillation can be performed on single hops as well as longer hops, as long as classical communication is available across those hops. When end-users place requests for entanglement, the network must execute a mechanism to deliver entanglement at the necessary rate and fidelity, besides other metrics such as latency. This must be done through a combination of good entanglement generation, efficient distillation and smart routing protocols. Hence, it is essential to develop scalable strategies for these processes that can meet the needs of different applications.
@@
In this paper, we focus specifically on entanglement distillation across single and multiple hops. Entanglement can be distilled using local operations and classical communication (LOCC), which can be one- or two-way [2]. Classic $2$-to-$1$ protocols require the nodes to perform two-way LOCC whereas protocols based on quantum error correction (QEC) only require one-way LOCC, significantly improving efficiency and reducing latency at the cost of increased complexity of LO. We consider QEC-based distillation protocols for second-generation quantum repeaters, where heralded entanglement generation addresses photon loss and QEC is used to correct errors introduced during LO [3]. Since recent years have witnessed an exciting amount of experimental demonstrations of QEC in different hardware technologies [4, 5, 6, 7, 8, 9, 10], these protocols could indeed be implemented in the next few years. Our goal here is to demonstrate a novel approach to adaptively improve end-to-end metrics that combine both rate and fidelity. Hence, to highlight the key insights and keep the setup simple, we incorporate noise in the channel but assume that all other components are noiseless. We emphasize that our innovation can easily be applied to realistic noise settings.
@@
We consider a linear chain of odd number of repeaters where each repeater generates entanglement with its neighboring nodes, distills, and then swaps them. The process continues until the end nodes have shared Bell pairs that they have distilled themselves. But independent of the length of the chain, we set the maximum number of distillation rounds to $3$. The entanglement generation is deterministic but the created state is a Werner state, where the Werner parameter could be interpreted as incorporating channel loss as well as noise in LO. These are the minimal assumptions we need to demonstrate our adaptive protocol. Conventional schemes apply the same distillation, or purification, procedure in each round to generate end-to-end entanglement of desired fidelity. However, we start with an $\llbracket n,k,d\rrbracket$ stabilizer code, $C_{1}$, with low rate, $R=\frac{k}{n}$, to correct many errors and improve the fidelity after the first round of distillation at single hops. After swapping the resulting $k$ Bell pairs, we adaptively switch to a different $\llbracket n,k^{\prime}>k,d^{\prime}\rrbracket$ code, $C_{2}$, that is able to further improve the fidelity while having a higher rate $R^{\prime}=\frac{k^{\prime}}{n}>R$. This is done carefully only when the input fidelity to the first round is just large enough to ensure that the resulting (output) fidelity is above the input fidelity threshold for the second code. For the third round of distillation, we adapt once again by picking yet another $\llbracket n,k^{\prime\prime}>k^{\prime}>k,d^{\prime\prime}\rrbracket$ code, $C_{3}$, if the channel operating point, i.e., depolarizing rate, is suitable. We fix $n$ throughout so that LO need not change greatly.
@@
Our simulations are conducted using the following codes: $C_{1}:\llbracket 9,1,3\rrbracket,C_{2}:\llbracket 9,2,3\rrbracket,C_{3}:\llbracket 9,3,3\rrbracket$. We compare four protocols: $\mathcal{P}_{1}:C_{1}-C_{1}-C_{1}$ that uses code $C_{1}$ all the time, whereas the other three, $\mathcal{P}_{2}:C_{1}-C_{2}-C_{2}$, $\mathcal{P}_{3}:C_{1}-C_{2}-C_{3}$, and $\mathcal{P}_{4}:C_{2}-C_{2}-C_{2}$, show the regions where they are advantageous. Our metric for comparison is a combination of overall rate (not just the code rate) and fidelity, which we call efficiency. Such adaptive QEC for entanglement distillation in quantum repeaters is unprecedented and opens up exciting new ways to achieve end-to-end metrics of quantum network applications. These can be easily modified for more realistic noise settings and the advantages will still qualitatively hold, but for a more sophisticated definition of networking operation point. We leave those deep investigations, especially a general construction of these code sequences as a function of target metrics, the associated fault tolerance requirements, and resource estimates, for future work.
@@
The paper is organized as follows. Section II introduces the necessary technical background, Section III discusses our adaptive protocol, Section IV demonstrates the simulation results for our protocol, and Section V concludes the paper by exploring avenues for future work.
@@ II-B Error Syndromes
Errors in a quantum system can be detected using the stabilizer formalism by measuring error syndromes. When an error $E$ acts on a code state |$\ket{\psi}$, it may commute or anti-commute with stabilizer operators $S_{i}$:
@@
This means that measuring the stabilizer will give an eigenvalue of $\lambda_{i}=\pm 1$, where $-1$s indicate the presence of an error. The error syndrome is a binary vector that records these eigenvalues: $\mathbf{s}=(\lambda_{1},\lambda_{2},\dots,\lambda_{m})$. For example, consider an error $E=XII$ acting on the $\llbracket 3,1,1\rrbracket$ code defined earlier. We calculate the syndrome as follows:
$ES_{1}\ket{\psi}$ $=XII\cdot ZZI\ket{\psi}=-ZZI\cdot XII\ket{\psi}=-S_{1}E\ket{\psi},$
$ES_{2}\ket{\psi}$ $=XII\cdot IZZ\ket{\psi}=IZZ\cdot XII\ket{\psi}=+S_{2}E\ket{\psi}.$
@@ II-C Depolarizing Noise Channel
where $p$ is the depolarizing probability. This means that with probability $1-p$, the qubit remains unchanged, and with probability $\frac{p}{3}$, one of the three Pauli errors $X,Y,Z$ is applied.
@@ II-E Distillable Entanglement
For a Werner state with fidelity $F$, the distillable entanglement obtained using one-way hashing is [11]:
@@
When we have three nodes, let $W_{1}$ be the parameter between node 1 and node 2, and $W_{2}$ be the parameter between node 2 and node 3. Then, we have the effective Werner parameter
@@ II-F Distillation based on Error Correction
The protocol begins with $n$ Bell States shared between Alice and Bob. Instead of sending them directly, Alice projects these states to the subspace of a stabilizer code. Since we use Bell states, the encoding can be implemented via stabilizer measurements on the initial Bell pairs. After encoding, Alice obtains $k$ encoded Bell pairs along with a set of stabilizers. The process is as follows [12, 13]:
@@ II-G Protocol Rate
A higher protocol rate indicates a more efficient high-fidelity entanglement distribution process.
@@ III-A Stabilizer Codes
In this work, we obtain stabilizer codes from the CodeTables database [14] and then transform them into standard form as outlined in [15, Chapter 10]. We describe the codes in the form of matrices whose entries are Pauli operators. Each row of the stabilizer matrix $H$ (resp. logical operator matrix $\overline{X}$ or $\overline{Z}$) corresponds to an independent stabilizer (resp. logical $X$ or logical $Z$) of the code. To illustrate the main ideas, we focus on the $\llbracket 9,1,3\rrbracket$, $\llbracket 9,2,3\rrbracket$ and $\llbracket 9,3,3\rrbracket$ codes shown in Table I. As an initial test for uncovering structural properties relevant to distillation, we also examine the well-known $\llbracket 5,1,3\rrbracket$ and $\llbracket 7,1,3\rrbracket$ codes. However, we emphasize that our method can be suitably adapted to other sequences of stabilizer codes.
@@ III-C Number of Distillation Rounds
In Fig. 3, the title of each subfigure follows the format “C-X-X", where “C" indicates that distillation is performed in that round, and “X" means that distillation is not performed in the corresponding round. Legends beginning with “0R" indicate cases without repeaters (i.e., these are just logical error rates for the codes). For example:
(a) represents performing only the first round of distillation. “0R, $\llbracket 5,1,3\rrbracket$" and “0R, $\llbracket 7,1,3\rrbracket$" denote directly establishing the entanglement link between Alice and Bob without repeaters, using the $\llbracket 5,1,3\rrbracket$ and $\llbracket 7,1,3\rrbracket$ codes, respectively. These curves are simply the logical error rates of the codes, plotted as input fidelity vs output fidelity. In this context, “$\llbracket 5,1,3\rrbracket$-X-X" indicates that only the first round of distillation is performed, using the $\llbracket 5,1,3\rrbracket$ code.
@@ III-E Assumptions
For our error correction-based distillation model, we use distillable entanglement as part of the end-to-end metric. To simplify the model, we assume that the only source of noise is the depolarizing channel, while all other components are ideal. This means that whenever we attempt to generate entanglement between two nodes, it will always succeed with $100\%$ probability, and there is no loss in the channel.
@@
The generated Bell pairs are considered ideal, and we simplify the noise model by assuming twirling is applied to produce depolarized Bell pairs. Quantum memories are assumed to hold qubits indefinitely without any decoherence or loss. Additionally, distillation and swapping operations are treated as ideal, with no errors.
These are the minimal assumptions that we need to demonstrate our adaptive error correction protocol. But we emphasize that our methods can be suitably adapted to realistic noise settings, and the advantages will qualitatively hold true.
@@ III-F Introduction to the Algorithm
Our simulation setup is built on MATLAB. When we try to build our algorithm, there are two major issues that need to be solved. The first one is the noise generation problem. When generating random Pauli noise, we expect it to be truly random. As is well known, random numbers generated by modern classical computers are all pseudo-random. Thus, generating results that approximate true randomness correctly is a significant goal for our algorithm. The second issue arises from the fact that the final result represents a probability distribution, which means we need a large number of noise simulations to obtain reliable final results. Based on our experiments, each mapping point between input fidelity $F_{\rm in}$ and output fidelity $F_{\rm out}$ requires at least 1,000,000 simulations to obtain a relatively accurate result. As a result, accelerating this process becomes the core goal of the algorithm.

### v2

# Adaptive Entanglement Distillation

@@ Abstract
Quantum network applications impose a variety of requirements on entanglement resources in terms of rate, fidelity, latency, and more. The repeaters in the quantum network must combine good methods for entanglement generation, effective entanglement distillation, and smart routing protocols to satisfy these application requirements. In this work, we focus on entanglement distillation in a linear chain of quantum repeaters. While conventional approaches reuse the same distillation scheme over multiple hop lengths after entanglement swaps, we propose a novel adaptive quantum error correction (QEC) scheme that boosts end-to-end metrics. Specifically, depending on the network operating point, we adapt the code used in distillation over successive rounds to monotonically increase the rate while also improving fidelity. We demonstrate the effectiveness of this strategy using three codes, with parameters $\llbracket 9,1,3\rrbracket$, $\llbracket 9,2,3\rrbracket$, $\llbracket 9,3,3\rrbracket$, and a new performance metric, efficiency, that incorporates both overall rate and fidelity. Since the minimum input fidelity for QEC-based distillation is high, we then extend our study to include non-QEC-based purification protocols, specifically DEJMPS since it outperforms others. We compare the performance of end-to-end DEJMPS against adapting from DEJMPS to QEC once DEJMPS improves the initial fidelity to the threshold for QEC. Through a refined efficiency metric, we illuminate the regime where QEC is beneficial. These results provide a detailed outlook for entanglement purification and distillation in first and second generation quantum repeaters.
@@ I Introduction
Quantum networks fundamentally differ from classical networks since they rely upon entanglement distribution across the network nodes. Since direct transmission of quantum information across long distances suffers from an exponential scaling of loss [Pirandola-natcomm17], quantum repeaters are critical for wide area quantum networks. A repeater generates entanglement across single hops – short links with its neighboring nodes, which could be end-users or other repeaters – and performs entanglement swaps to create longer-range entanglement across longer hops. Since each link introduces loss, and swaps involve local operations that are noisy, entanglement distillation is used to distill fewer high fidelity entangled states from many noisy entangled states. Such distillation can be performed on single hops as well as longer hops, as long as classical communication is available across those hops. When end-users place requests for entanglement, the network must execute a mechanism to deliver entanglement at the necessary rate and fidelity, besides other metrics such as latency. This must be done through a combination of good entanglement generation, efficient distillation and smart routing protocols. Hence, it is essential to develop scalable strategies for these processes that can meet the needs of different applications.
@@
In this paper, we focus specifically on entanglement distillation across single and multiple hops. Entanglement can be distilled using local operations and classical communication (LOCC), which can be one- or two-way [Bennett-pra96]. Classic 22-to-11 protocols require the nodes to perform two-way LOCC whereas protocols based on quantum error correction (QEC) only require one-way LOCC, significantly improving efficiency and reducing latency at the cost of increased complexity of LO. We consider QEC-based distillation protocols for second-generation quantum repeaters, where heralded entanglement generation addresses photon loss and QEC is used to correct errors introduced during LO [Muralidharan-scirep16]. Since recent years have witnessed an exciting amount of experimental demonstrations of QEC in different hardware technologies [Berthusen-arxiv24, Bluvstein-nature24, daSilva-arxiv24, Gupta-nature24, Lescanne-natphys20, Reichardt-arxiv24, SalesRodriguez-arxiv24], these protocols could indeed be implemented in the next few years. Our goal here is to demonstrate a novel approach to adaptively improve end-to-end metrics that combine both rate and fidelity. Hence, to highlight the key insights and keep the setup simple, we incorporate noise in the initial generation of Bell pairs but assume that all other components are noiseless. In principle, memory decoherence or noise in LO can be mapped back to the initial fidelity of the Bell pairs, potentially with some Pauli twirling. We emphasize that our innovation can be adapted to realistic noise settings with careful considerations of decoherence and fault tolerance. Since these considerations will distract from the central focus of this work, and warrant an independent investigation on their own, we leave them for subsequent follow-up work.
@@
We consider a linear chain of odd number of repeaters where each repeater generates entanglement with its neighboring nodes, distills, and then swaps them. The process continues until the end nodes have shared Bell pairs that they have distilled themselves. But independent of the length of the chain, we set the maximum number of distillation rounds to $3$, as explained later. The entanglement generation is deterministic but the created state is a Werner state, where the Werner parameter could be interpreted as incorporating decoherence as well as noise in LO. These are the minimal assumptions we need to demonstrate our adaptive protocol. Conventional schemes apply the same distillation, or purification, procedure in each round to generate end-to-end entanglement of desired fidelity. However, we start with an $\llbracket n,k,d\rrbracket$ stabilizer code, $C_{1}$, with low rate, $R=\frac{k}{n}$, to correct many errors and improve the fidelity after the first round of distillation at single hops. After swapping the resulting $k$ Bell pairs, we adaptively switch to a different $\llbracket n,k^{\prime}>k,d^{\prime}\rrbracket$ code, $C_{2}$, that is able to further improve the fidelity while having a higher rate $R^{\prime}=\frac{k^{\prime}}{n}>R$. This is done carefully only when the input fidelity to the first round is just large enough to ensure that the resulting (output) fidelity is above the input fidelity threshold for the second code. For the third round of distillation, we adapt once again by picking yet another $\llbracket n,k^{\prime\prime}>k^{\prime}>k,d^{\prime\prime}\rrbracket$ code, $C_{3}$, if the channel operating point, i.e., depolarizing rate, is suitable. We fix $n$ throughout so that LO need not change greatly.
@@
For illustration of the idea, our simulations are conducted using the following codes: $C_{1}:\llbracket 9,1,3\rrbracket,C_{2}:\llbracket 9,2,3\rrbracket,C_{3}:\llbracket 9,3,3\rrbracket$. We compare four protocols, $\mathcal{P}_{1}:C_{1}-C_{1}-C_{1}$, $\mathcal{P}_{2}:C_{1}-C_{2}-C_{2}$, $\mathcal{P}_{3}:C_{1}-C_{2}-C_{3}$, and $\mathcal{P}_{4}:C_{2}-C_{2}-C_{2}$, and show the regions where each one is advantageous. Our metric for comparison is a combination of overall rate (not just the code rate) and fidelity, which we call efficiency. Such adaptive QEC for entanglement distillation in quantum repeaters is unprecedented and opens up exciting new ways to achieve end-to-end metrics of quantum network applications. These can be adapted for more realistic noise settings and the advantages will still qualitatively hold, but for a more sophisticated definition of network operating point. We leave those deep investigations, e.g., a general construction of code sequences as a function of target metrics, their fault tolerance requirements, and resource estimates, for future work.
In the above exploration of adaptive QEC, the minimum required input fidelities for the codes are high, $\gtrapprox 90\%$. For near-term implementation, it is important to keep the codes small but plan for low input fidelities, $\approx 70-80\%$, especially while accounting for decoherence and noise in LO. Even in the long-term, the fault-tolerant thresholds for large quantum low-density parity-check (QLDPC) codes are also expected to be high. Hence, we consider the use of classic non-QEC-based entanglement purification protocols such as BBPSSW [PhysRevLett.76.722_BBPSSW] and DEJMPS [PhysRevLett.77.2818_DEJMPS] to improve initial fidelities to the threshold of the above QEC codes [^1] Note that in the QEC literature this is called a ‘pseudo-threshold’ since it is for a specific code and not a property of a family of codes. A ‘threshold’ usually refers to a family. Here, we will use ‘threshold’ for convenience.. We first show that DEJMPS with no twirling comprehensively outperforms BBPSSW (with/without twirling), so we restrict our consideration to only DEJMPS for the rest of the paper. We illuminate the asymmetric nature of the Pauli error probabilities for DEJMPS and show that it enables the protocol to rapidly improve fidelities in each round. Then we use DEJMPS to improve initial fidelities to the threshold of the $\llbracket 9,3,3\rrbracket$ code, chosen for its high rate compared to the other two codes. We highlight the emergence of “checkpoints” in the input-output fidelity plot as the minimum required number of DEJMPS rounds decreases with increasing input fidelity. After DEJMPS improves the input fidelity above the code threshold, we consider switching to one round of the QEC protocol and compare it to continuing with further DEJMPS rounds to achieve the same performance. We refine our previously defined efficiency metric to account for the subtleties of this setting. Using this refined metric, we highlight the regions where pure DEJMPS outperforms the switch to QEC and vice-versa. Importantly, we observe that it is the high rate of the QEC code that provides an edge at very high input fidelities, rather than the output fidelity.
This suggests that the design of second generation repeaters must carefully inspect the choice and value of a QEC code to replace or co-exist with simpler protocols such as DEJMPS. We expect that QEC would become a preferable choice under a fault-tolerant setting, but the details of the code and protocol will be critical to ensure superior performance to DEJMPS. We are not aware of prior works that have closely examined the evolution from first- to second-generation repeaters, so we think that our initial investigation will lead to many more careful studies of this important technological aspect.
@@
The paper is organized as follows. Section II introduces the necessary technical background, Section III discusses our adaptive protocol, Section IV demonstrates the simulation results for our protocol, Section V explores the extension to non-QEC-based protocols, and Section VI concludes the paper by exploring avenues for future work. Appendix A provides analytical details of the comparison between BBPSSW and DEJMPS in Section V.
@@ II-B Error Syndromes
[…]
This means that measuring the stabilizer will give an eigenvalue of $\lambda_{i}=\pm 1$, where 1-1s indicate the presence of an error. The error syndrome is a binary vector that records these eigenvalues: $\mathbf{s}=(\lambda_{1},\lambda_{2},\dots,\lambda_{m})$. For example, consider an error $E=XII$ acting on the $\llbracket 3,1,1\rrbracket$ code defined earlier. We calculate the syndrome as follows:
[…]
@@ II-C Depolarizing Noise Channel
where $p$ is the depolarizing probability. This means that with probability 1p1-p, the qubit remains unchanged, and with probability $\frac{p}{3}$, one of the three Pauli errors $X,Y,Z$ is applied.
@@ II-E Distillable Entanglement
For a Werner state with fidelity $F$, the distillable entanglement obtained using one-way hashing is [Patil2024EntanglementRouting]:
@@
When we have three nodes, let $W_{1}$ be the parameter between nodes 1 and 2, and $W_{2}$ be the parameter between nodes 2 and 3. Then, after a swap at node 2, we have the effective Werner parameter between nodes 1 and 3,
@@ II-F Distillation based on Error Correction
The protocol begins with $n$ Bell States shared between Alice and Bob. Instead of sending them directly, Alice projects these states to the subspace of a stabilizer code. Since we use Bell states, the encoding can be implemented via stabilizer measurements on the initial Bell pairs. After encoding, Alice obtains $k$ encoded Bell pairs along with a set of stabilizers. The process is as follows [Wilde-isit10, Rengaswamy-quantum24]:
@@ II-G Protocol Rate
A higher protocol rate indicates a more efficient high-fidelity entanglement distribution process. Later, we will define an efficiency metric that combines this rate with fidelity.
@@ III-A Stabilizer Codes
In this work, we obtain stabilizer codes from the CodeTables database [Grassl:codetables] and then transform them into standard form as outlined in [NielsenChuang2010, Chapter 10]. We describe the codes in the form of matrices whose entries are Pauli operators. Each row of the stabilizer matrix $H$ (resp. logical operator matrix $\overline{X}$ or $\overline{Z}$) corresponds to an independent stabilizer (resp. logical $X$ or logical $Z$) of the code. To illustrate the main ideas, we focus on the $\llbracket 9,1,3\rrbracket$, $\llbracket 9,2,3\rrbracket$ and $\llbracket 9,3,3\rrbracket$ codes shown in Table I. As an initial test for uncovering structural properties relevant to distillation, we also examine the well-known $\llbracket 5,1,3\rrbracket$ and $\llbracket 7,1,3\rrbracket$ codes. However, we emphasize that our method can be suitably adapted to other sequences of stabilizer codes.
@@ III-C Number of Distillation Rounds
In Fig. 2, the title of each subfigure follows the format “C-X-X”, where “C” indicates that distillation is performed in that round using either code, and “X” means that distillation is not performed in the corresponding round. Legends beginning with “0R” indicate cases without repeaters (i.e., these are just logical error rates for the codes). For example:
(a) represents performing only the first round of distillation. “0R, $\llbracket 5,1,3\rrbracket$” and “0R, $\llbracket 7,1,3\rrbracket$” denote directly establishing the entanglement link between Alice and Bob without repeaters, using the $\llbracket 5,1,3\rrbracket$ and $\llbracket 7,1,3\rrbracket$ codes, respectively. These curves are simply the logical error rates of the codes, plotted as input fidelity vs output fidelity. In this context, $\llbracket 5,1,3\rrbracket$-X-X” indicates that only the first round of distillation is performed, using the $\llbracket 5,1,3\rrbracket$ code.
@@ III-E Assumptions
For our error correction-based distillation model, we use distillable entanglement as part of the end-to-end metric. To simplify the model, we assume that the only source of noise is the depolarizing channel, while all other components are ideal. This means that whenever we attempt to generate entanglement between two nodes, it will always succeed with $100\%$ probability, and there is no loss in the channel. Our primary goal here is to demonstrate the effectiveness of our adaptive distillation protocol in the simplest nontrivial setting.
@@
The generated Bell pairs are considered ideal, and we simplify the noise model by assuming twirling is applied to produce depolarized Bell pairs. Quantum memories are assumed to hold qubits indefinitely without any decoherence or loss. Note that they only need to hold qubits until they are consumed in the next swap or distillation, which will not be very long. Additionally, distillation and swapping operations are treated as ideal, with no errors. If we consider circuit-level errors, then the entire scheme must be made fault-tolerant for our sequence of codes, which warrants an entire paper for itself that considers all the subtle details. We do intend to perform that analysis as a natural follow-up to this work.
@@ III-F Introduction to the Algorithm
Our simulation setup is built on MATLAB. When we try to build our algorithm, there are two major issues that need to be solved. The first one is the noise generation problem. When generating random Pauli noise, we expect it to be truly random. As is well-known, random numbers generated by modern classical computers are all pseudo-random. Thus, generating results that approximate true randomness correctly is a significant goal for our algorithm. The second issue arises from the fact that the final result represents a probability distribution, which means we need a large number of noise simulations to obtain reliable final results. Based on our experiments, each mapping point between input fidelity $F_{\rm in}$ and output fidelity $F_{\rm out}$ requires at least 1,000,000 simulations to obtain a relatively accurate result. As a result, accelerating this process becomes the core goal of the algorithm.